OpenStudy (anonymous):

x/(x - 3) less than or equal to -8/(x - 6) Solve for x using any method

OpenStudy (anonymous):

\[x \div(x - 3) \le -8/(x - 6)\]

OpenStudy (anonymous):

I got to [(x - 1)(x + 2)]/[(x - 3)(x - 6) le 0 and dunno where to go from there.

OpenStudy (anonymous):

I wish I could rage quit this class lol.

OpenStudy (anonymous):

Your step wasn't exactly right. Look there: http://www.wolframalpha.com/input/?i=x%C3%B7%28x%E2%88%923%29%E2%89%A4%E2%88%928%2F%28x%E2%88%926%29

jhonyy9 (jhonyy9):

before all you need push the conditions that x not can be equal 3 and 6 becouse than those fractions are undefined so x/(x-3) +8/(x-6) <= 0 x(x-6)+8(x-3)<=0 x2-6x+8x-24<=0 x2 +2x -24<=0

jhonyy9 (jhonyy9):

x2 +2x -24=(x+6)(x-4)

jhonyy9 (jhonyy9):

so than x+6=0 so x=-6 and x-4=0 so x=4 those are zeros for this function

OpenStudy (anonymous):

I'm trying to bring this all in lol

jhonyy9 (jhonyy9):

now you need check for what intervale(s) is less than zero - becous you know that for -6 and 4 is equal zero

jhonyy9 (jhonyy9):

hope that was understandably

jhonyy9 (jhonyy9):

???

OpenStudy (anonymous):

Gimme a sec

OpenStudy (anonymous):

How do I check for what you said to do with the intervals??

jhonyy9 (jhonyy9):

ok try

OpenStudy (anonymous):

Try what? I'm confused!

jhonyy9 (jhonyy9):

why ? you know that for 3 and 6 those fractions are undefined and for -6 and 4 are zeros

OpenStudy (anonymous):

Yeah so how do I solve for x from there?

jhonyy9 (jhonyy9):

now you need get in what intervale are less than zero

jhonyy9 (jhonyy9):

so one function when is less than zero ?

jhonyy9 (jhonyy9):

you know that just when for this value of x the resulted number is negativ - right ???

OpenStudy (anonymous):

No I don't understand, I'm very sorry.

jhonyy9 (jhonyy9):

when one number is less than zero ?

jhonyy9 (jhonyy9):

comen i like to learned it

jhonyy9 (jhonyy9):

learne it sorry

jhonyy9 (jhonyy9):

what is your answer ?

jhonyy9 (jhonyy9):

sorry i need to go

OpenStudy (anonymous):

Idk man I give up I have no idea what you're trying to tell me

jhonyy9 (jhonyy9):

why this is whery easy

jhonyy9 (jhonyy9):

your exercise what youn have wrote one function when is less or equal zero

jhonyy9 (jhonyy9):

becouse there is x/(x-3)+8/(x-6) <=0 after you add to both sides 8/(x-6)

OpenStudy (anonymous):

I understand that the zeros are 4 and -6 and x can't = 3 or 6. I just dunno what the asnwer is from there

OpenStudy (anonymous):

Fraction is <=0 when: 1. numerator is <=0 and denominator is >0 OR 2. numerator is >=0 and denominator is <0

jhonyy9 (jhonyy9):

now you need to check that for what intervale is less than zero ?

jhonyy9 (jhonyy9):

probe you understanding right what have wrote to you marina

OpenStudy (anonymous):

I still don't get it lol omg this is so lame I give up.

OpenStudy (anonymous):

I dont understand how to find the interval at all. This is going right over my head

jhonyy9 (jhonyy9):

Marina can you help this person ?

jhonyy9 (jhonyy9):

because i need to going now

OpenStudy (anonymous):